OpenStudy (anonymous):

A family has three children. Assuming the chance of either a boy or a girl is equal (50% each), what is the chance that the family has three girls given that the first child is a girl.

OpenStudy (anonymous):

\[\frac{1}{4}\]

OpenStudy (anonymous):

you know that the first child is already a girl so the question really is 'what it the probability they have two more girls?' assuming independence you just multiply the probabilities together \[\frac{1}{2} \times \frac{1}{2}\]